OpenStudy (anonymous):

Given: m∠1+m∠2=100 and m∠1=80 Prove: m∠2=20

OpenStudy (anonymous):

http://tinypic.com/r/2yw8h9t/6 here is the problem

OpenStudy (mayankdevnani):

m<1+m<2=100 80+m<2=100 m<2=100-80=20

OpenStudy (mayankdevnani):

ok @ElizabethScarlette

OpenStudy (mayankdevnani):

@thivitaa am i right?

OpenStudy (anonymous):

1. Given 2. Substitution 3. Subtraction Property of Equality

OpenStudy (anonymous):

in that case the 3rd and has to be an obtuse and le to add the 3rd angle then subtract by the 2nd angle to get the sum of that